# Galactic Archaeology with Gaia

### v1

@@ 2.1 Gaia Data Release 1
On September 14 2016 the very first Gaia Data Release was delivered to the eager, but somewhat unready, community to forever change the way Galactic Astronomy is done. However, the unprecedented all-sky $\sim$1 billion-strong stars only came with mean positional coordinates, RA and Dec — in Gaia DR1, proper motions, and parallaxes were limited to a tiny $\sim$ 2 million subset of the brightest stars that had been seen previously by Hipparcos (Gaia Collaboration et al., 2016a). The size of this so-called Tycho-Gaia Astrometric Solution (TGAS, Lindegren et al., 2016) was unfortunately too small to explore in detail the behaviour of the stellar halo in the low-metallicity regime. Surprisingly, however, a significant number of metal-rich stars ([Fe/H]$>-1$) on halo-like orbits were detected in the Solar neighborhood by Bonaca et al. (2017). These metal-rich stars comprised approximately half of the halo sample near the Sun and exhibited a clear prograde bias compared to their much more isotropic low-metallicity counterparts. Using measurements from the APOGEE survey, this metal-rich population is shown to follow the sequence of the old Milky Way disc in the $\alpha$-[Fe/H] plane (see Fig. 3). Weighing up the observational evidence and comparing to the numerical simulations, Bonaca et al. (2017) concluded that this metal-rich halo population must be of in-situ origin and suggested the disc heating by merger events as the key pathway to its formation. While not explicitly connecting the in-situ halo to the last massive merger, the inventive and dexterous use of the limited DR1 data and the insightful comparison to the available simulations makes the pioneering paper by Bonaca et al. (2017) a clear highlight of the early Gaia exploration which helped to set up the stage for the extensive DR2 analysis.
@@
Aside from the TGAS, the rest of the billion stars in Gaia DR1 would have their detailed astrometry published later, as part of the DR2. Not prepared to wait, several groups decided to combine Gaia DR1 with earlier astrometric data from elsewhere thus taking advantage of the enormous, uniform, and high-quality dataset Gaia was providing, as well as the large temporal baseline between Gaia and other surveys. At least three such catalogues were made: the first two obtained proper motion estimates simply by cross-matching Gaia DR1 with the data from PPMXL [^1] PPMXL itself is based on 2MASS and USNO. (HSOY, see Altmann et al., 2017), and PS1 (Tian et al., 2017). The third proper motion catalogue was made by Sergey Koposov following a rather different strategy to connect Gaia DR1 and SDSS. Koposov realised that the Gaia data could be used to fix the systematic biases in the astrometric calibration of the SDSS itself before using the two to measure positional offsets. Unfortunately, the details of the SDSS re-calibration and the catalogue construction were not described in a stand-alone paper, but the tests of the catalogue performance were published in Deason et al. (2017b) and de Boer et al. (2018).
@@
Belokurov et al. (2018b) elucidate the genesis of the radially-anisotropic halo component with the analysis of high-resolution zoom-in simulations of Milky Way formation. They show that in the mock stellar halo samples at host-centric distances similar to Solar, the orbital anisotropy is a strong function of the progenitor’s mass and its accretion time in agreement with earlier studies (see e.g Deason et al., 2013a). Across a wide range of accretion events explored, only the most massive progenitors with dark matter masses of the order of $10^{11}M_{\odot}$ can yield high anisotropy $\beta>0.8$ provided the mergers took place at $21<z<$ 2. In their simulation suite, this period in the Galaxy’s life also coincided with the phase of active disc growth. The growing disc seemed to help make the orbit of the in-falling massive satellite more eccentric. Most importantly, however, there is a conspicuous convergence in that many independent lines of evidence point to a particular accretion history for our Galaxy. Namely, the one in which a single massive dwarf galaxy was accreted 8-11 Gyr ago, as previously hypothesized by Deason et al. (2013a). As predicted by Amorisco (2017), the orbit of the dwarf radialised during its interaction with the Milky Way, populating the inner halo with stellar tidal debris on nearly radial orbits. The resulting high eccentricity of the main stellar halo component also explains the low net spin measured by Deason et al. (2017b). Finally, the relatively high metallicity of the eccentric halo stars in the Solar neighbourhood agrees with the expectation for stellar contents of a massive dwarf galaxy in accordance with the established mass-metallicity relation (see e.g. Kirby et al., 2013). Belokurov et al. (2018b) conclude that the sausage-like feature in the local phase-space ought to be the result of an ancient accretion of a single, massive satellite.
@@
Phase-space analysis of the kind described above is appropriate for data samples limited to the Solar neighborhood. It is also very convenient as it does not require an assumption of the gravitational potential of the Galaxy and can be re-produced easily, including in applications to mock data from numerical simulations. However, as demonstrated by e.g. Helmi and de Zeeuw (2000), the identification of merger debris is more straightforward and more efficient in the integrals-of-motion space. This is exactly what is attempted in Helmi et al. (2017) where a low-metallicity halo sample is selected using a cross-match between TGAS and RAVE (see Steinmetz et al., 2006). Unfortunately, the resulting number of stars ($\sim 1000$) is too low to discern clearly any global patterns, although the presence of a conspicuous retrograde population is spotted. Instead, similarly to Deason et al. (2017b) and Belokurov et al. (2018b), Myeong et al. (2018b) used the much larger SDSS-Gaia sample which contained $>10,000$ halo stars across a wide range of metallicity. They scrutinize the behaviour of the halo in the action space in different [Fe/H] bins and detect the prevalence of highly eccentric orbits at $-2<$[Fe/H]$<-1$, for which they agree the simplest explanation is the in-fall of one satellite.
@@
Figure 6: Left: HRD for stars with high tangential velocities (typical halo kinematics) in Gaia DR2. Note the split main sequence. [Adapted from Gaia Collaboration et al. (2018a)]. Middle: High tangential velocity stars separated into blue and red populations with a model isochrone corresponding to [Fe/H]$=-0.75$ and age of 11.5 Gyr. Right: Toomre diagram (also see Fig. 3) for blue and red stars selected in the previous panel. [Adapted from Haywood et al. (2018)].
@@ 2.2 Gaia Data Release 2
    The motion of the stars with [Fe/H]$<-1$ within $\sim 20-30$ kpc appears synchronised, which in turn explains the global spatial structure of the halo (Deason et al., 2018; Iorio and Belokurov, 2019)
@@
HRD dichotomy. The striking apparent split in the Hertzsprung-Russell diagram (HRD) of the nearby stars with high tangential velocity — i.e. on halo-like orbits — was first presented and discussed in one of the Gaia DR2 science demonstration papers (Gaia Collaboration et al., 2018a, see Fig. 6). The authors notice two separate HRD sequences — blue and red according to their $G_{BP}-G_{RP}$ colour — most visibly disconnected around the main sequence turn-off region, which they analyze and interpret to be produced by two halo populations with distinct mean metallicities, [Fe/H]$\sim-1.3$ and [Fe/H]$\sim-0.5$. Gaia Collaboration et al. (2018a) report clear similarities with the bi-modality in the metallicity distribution of the Galactic GCs, which is known to arise due to contributions from clusters predominantly accreted from dwarf galaxies and those formed in situ in the Milky Way. They speculate that the two halo populations must have followed distinct formation channels and surmise that there may be a connection between the two halo HRD sequences and the chemical sequences highlighted earlier in Nissen and Schuster (2010) and Bonaca et al. (2017). Haywood et al. (2018) take a deeper look at the Gaia DR2 HRD to solve the puzzle of the dichotomy of the stellar halo. They demonstrate that while both populations have orbits distinct from the disc, there is a clear difference in the kinematics of the stars in the blue and red sequences. Haywood et al. (2018) show that the red sequence is composed of stars on mostly prograde orbits with lower eccentricities and lower total energies. The blue sequence stars have little angular momentum, have higher eccentricities, and reach out further into the halo. The HRD sequences indeed correspond almost perfectly to the two chemical sequences found by Nissen and Schuster (2010): the red one contains stars with higher $\alpha$-abundances and the blue one with lower $\alpha$-abundances. The connection between kinematics and detailed chemistry is also made in Helmi et al. (2018), who together with Haywood et al. (2018) appear inspired by the earlier analysis of the halo chemistry using APOGEE (see Hawkins et al., 2015; Fernández-Alvar et al., 2018; Hayes et al., 2018). Haywood et al. (2018) associate the red sequence to the thick disc in agreement with Bonaca et al. (2017) and the blue sequence to the ancient massive merger as proposed in e.g. Deason et al. (2013a) and Belokurov et al. (2018b). Haywood et al. (2018) argue that the tightness of the blue sequence in the Gaia HRD is an argument in favour of the single dominant accretion event because contributions from many dwarf progenitors would otherwise bloat it.
@@
Orbital synchronicity. The use of the GCs as tracers highlighted one particular difficulty of the stellar halo studies with Gaia. Put simply, given the limited reach of the Gaia DR2 parallaxes, the vast extent of the halo appeared inaccessible. These observational hindrances can be circumvented using tools of Galactic Dynamics with which one can predict the overall density distribution of the halo using only a small sample of tracer kinematics constrained to the Solar neighborhood (see e.g. Eggen et al., 1962; Chiba and Beers, 2000). This is because the spatial appearance of the Galaxy is nothing but the superposition of the orbits of the individual stars and the density in the configuration space is linked to the kinematics through the continuity of the phase-space flow (Binney and Tremaine, 2008). In a similar vein but in a much more direct experiment, Deason et al. (2018) compute orbits of the metal-rich halo stars in the Solar vicinity and demonstrate that their apo-centres pile-up around Galactocentric distance $r\approx 20$ kpc (see Fig. 8). The location of this pile-up matches the break radius in the stellar halo density (see Section 1.1.2). As a sanity check, they also compare apo-centres of nearby main sequence stars to the apo-centres of the distant BHB stars that were originally used to detect the break and show that stars in the two populations, distant and local, turn around on their orbits at approximately the same distance. Thus three different tracer populations have been shown to have the same orbital behaviour: the nearby main sequence stars, the distant BHBs, and the GCs. Deason et al. (2018) demonstrate in a very candid way that the local metal-rich and radially anisotropic population is only a glimpse of a vast and dominant cloud of stellar debris that dictates the global properties of the stellar halo (e.g. the density break). They also provide the first constraint on the metallicity distribution function of the high-eccentricity population and show that its incidence drops by a factor of $\sim 2$ going from [Fe/H]$=-1$ to [Fe/H]$=-2$.
@@
Helmi et al. (2018) show that on the HRD, these stars are largely confined by isochrones with $-1.3<$[Fe/H]$<-0.9$ and ages of 10-13 Gyr. Helmi et al. (2018) cross-match their selected sample to APOGEE and see two narrow sequences in the $\alpha$-[Fe/H] plane, one mostly flat with $-1<$[Fe/H]$<-0.3$ and [$\alpha$/Fe]$\approx 0.25$ and one sloping down from [Fe/H]$\approx-1.3$, [$\alpha$/Fe] =0.25$\approx$=0.25≈ = 0.25 to [Fe/H]$\approx-0.7$, [$\alpha$/Fe] $\approx 0.1$. Following Bonaca et al. (2017), they assign the high-$\alpha$ sequence to the in-situ population (see also Nissen and Schuster, 2010). Given that the two sequences co-exist over a range of metallicities, Helmi et al. (2018) conclude that the halo stars in the lower $\alpha$-[Fe/H] sequence must have formed in a separate, smaller galaxy. In view of the large spread of metallicity and the presence of low-$\alpha$ high-[Fe/H] stars this dwarf could not be much smaller than the LMC. Their interpretation of the two chemical sequences builds on the analysis of the halo stars in APOGEE by Hawkins et al. (2015), Fernández-Alvar et al. (2018) and Hayes et al. (2018). These earlier works provide a detailed chemo-dynamical comparison of the high- and low-$\alpha$ sequences and agree on the in-situ nature of the former. Crucially, however, in the absence of the Gaia data, these authors refrain from making a claim as to the exact nature of the accreted population, although Fernández-Alvar et al. (2018) report a significantly lower star-formation rate for the low-$\alpha$ population. Helmi et al. (2018)’s conclusion is not solely based on the accumulation of the observational evidence available at the time. Their insight is also the result of an earlier numerical exploration, a series of experiments simulating high mass-ratio mergers between the Milky Way and a dwarf satellite (Villalobos and Helmi, 2008). Helmi et al. (2018) find striking similarities between the Gaia DR2 data and one of the pre-existing simulations (once the simulation is appropriately re-scaled). Helmi et al. (2018) propose to give a name to the progenitor system (and presumably its tidal debris): Gaia-Enceladus.
Following the two papers that first identified the metal-rich highly-eccentric accreted halo population and proposed that it must be produced in a single massive accretion event, the vast tidal debris cloud is referred to as Gaia Sausage-Enceladus (GS/E).

### v2

@@ 2.1 Gaia Data Release 1
On September 14 2016 the very first Gaia Data Release was delivered to the eager, but somewhat unready, community to forever change the way Galactic Astronomy is done. However, the unprecedented all-sky $\sim$1 billion-strong stars only came with mean positional coordinates, RA and Dec — in Gaia DR1, proper motions, and parallaxes were limited to a tiny $\sim$ 2 million subset of the brightest stars that had been seen previously by Hipparcos (Gaia Collaboration et al., 2016a). The size of this so-called Tycho-Gaia Astrometric Solution (TGAS, Lindegren et al., 2016) was unfortunately too small to explore in detail the behaviour of the stellar halo in the low-metallicity regime. Surprisingly, however, a significant number of metal-rich stars ([Fe/H]$>-1$) on halo-like orbits were detected in the Solar neighourhood by Bonaca et al. (2017). These metal-rich stars comprised approximately half of the halo sample near the Sun and exhibited a clear prograde bias compared to their much more isotropic low-metallicity counterparts. Using measurements from the APOGEE survey, this metal-rich population is shown to follow the sequence of the old Milky Way disc in the $\alpha$-[Fe/H] plane (see Fig. 3). Weighing up the observational evidence and comparing to the numerical simulations, Bonaca et al. (2017) concluded that this metal-rich halo population must be of in-situ origin and suggested the disc heating by merger events as the key pathway to its formation. While not explicitly connecting the in-situ halo to the last massive merger, the inventive and dexterous use of the limited DR1 data and the insightful comparison to the available simulations makes the pioneering paper by Bonaca et al. (2017) a clear highlight of the early Gaia exploration which helped to set up the stage for the extensive DR2 analysis.
@@
Aside from the TGAS, the rest of the billion stars in Gaia DR1 would have their detailed astrometry published later, as part of the DR2. Not prepared to wait, several groups decided to combine Gaia DR1 with earlier astrometric data from elsewhere thus taking advantage of the enormous, uniform, and high-quality dataset Gaia was providing, as well as the large temporal baseline between Gaia and other surveys. At least three such catalogues were made: the first two obtained proper motion estimates simply by cross-matching Gaia DR1 with the data from PPMXL [^1] PPMXL itself is based on 2MASS and USNO. (HSOY, see Altmann et al., 2017), and PS1 (Tian et al., 2017). The third proper motion catalogue was made by Sergey Koposov following a rather different strategy to connect Gaia DR1 and SDSS (see Koposov, 2024, for details and the catalogue download). Koposov realized that the Gaia data could be used to fix the systematic biases in the astrometric calibration of the SDSS itself before using the two to measure positional offsets. Unfortunately, the details of the SDSS re-calibration and the catalogue construction were not described in a stand-alone paper, but the tests of the catalogue performance were published in Deason et al. (2017b) and de Boer et al. (2018).
@@
Belokurov et al. (2018b) elucidate the genesis of the radially-anisotropic halo component with the analysis of high-resolution zoom-in simulations of Milky Way formation. They show that in the mock stellar halo samples at host-centric distances similar to Solar, the orbital anisotropy is a strong function of the progenitor’s mass and its accretion time in agreement with earlier studies (see e.g Deason et al., 2013a). Across a wide range of accretion events explored, only the most massive progenitors with dark matter masses of the order of $10^{11}M_{\odot}$ can yield high anisotropy $\beta>0.8$ provided the mergers took place at $21<z<$ 2. In their simulation suite, this period in the Galaxy’s life also coincided with the phase of active disc growth. The growing disc seemed to help make the orbit of the in-falling massive satellite more eccentric. Most importantly, however, there is a conspicuous convergence in that many independent lines of evidence point to a particular accretion history for our Galaxy. Namely, the one in which a single massive dwarf galaxy was accreted 8-11 Gyr ago, as previously hypothesized by Deason et al. (2013a). As predicted by Amorisco (2017), the orbit of the dwarf radialized during its interaction with the Milky Way, populating the inner halo with stellar tidal debris on nearly radial orbits. The resulting high eccentricity of the main stellar halo component also explains the low net spin measured by Deason et al. (2017b). Finally, the relatively high metallicity of the eccentric halo stars in the Solar neighbourhood agrees with the expectation for stellar contents of a massive dwarf galaxy in accordance with the established mass-metallicity relation (see e.g. Kirby et al., 2013). Belokurov et al. (2018b) conclude that the sausage-like feature in the local phase-space ought to be the result of an ancient accretion of a single, massive satellite.
@@
Phase-space analysis of the kind described above is appropriate for data samples limited to the Solar neighourhood. It is also very convenient as it does not require an assumption of the gravitational potential of the Galaxy and can be re-produced easily, including in applications to mock data from numerical simulations. However, as demonstrated by e.g. Helmi and de Zeeuw (2000), the identification of merger debris is more straightforward and more efficient in the integrals-of-motion space. This is exactly what is attempted in Helmi et al. (2017) where a low-metallicity halo sample is selected using a cross-match between TGAS and RAVE (see Steinmetz et al., 2006). Unfortunately, the resulting number of stars ($\sim 1000$) is too low to discern clearly any global patterns, although the presence of a conspicuous retrograde population is spotted. Instead, similarly to Deason et al. (2017b) and Belokurov et al. (2018b), Myeong et al. (2018b) used the much larger SDSS-Gaia sample which contained $>10,000$ halo stars across a wide range of metallicity. They scrutinize the behaviour of the halo in the action space in different [Fe/H] bins and detect the prevalence of highly eccentric orbits at $-2<$[Fe/H]$<-1$, for which they agree the simplest explanation is the in-fall of one satellite.
@@
Figure 6: Left: HRD for stars with high tangential velocities (typical halo kinematics) in Gaia DR2 (Gaia Collaboration et al., 2018a). Note the split main sequence. Middle: High tangential velocity stars separated into blue and red populations with a model isochrone corresponding to [Fe/H]$=-0.75$ and age of 11.5 Gyr (Haywood et al., 2018). Right: Toomre diagram (also see Fig. 3) for blue and red stars selected in the previous panel. [Adapted from Gaia Collaboration et al. (2018a); Haywood et al. (2018)].
@@ 2.2 Gaia Data Release 2
    The motion of the stars with [Fe/H]$<-1$ within $\sim 20-30$ kpc appears synchronized, which in turn explains the global spatial structure of the halo (Deason et al., 2018; Iorio and Belokurov, 2019)
@@
HRD dichotomy. The striking apparent split in the Hertzsprung-Russell diagram (HRD) of the nearby stars with high tangential velocity — i.e. on halo-like orbits — was first presented and discussed in one of the Gaia DR2 science demonstration papers (Gaia Collaboration et al., 2018a, see Fig. 6). The authors notice two separate HRD sequences — blue and red according to their $G_{BP}-G_{RP}$ colour — most visibly disconnected around the main sequence turn-off region, which they analyse and interpret to be produced by two halo populations with distinct mean metallicities, [Fe/H]$\sim-1.3$ and [Fe/H]$\sim-0.5$. Gaia Collaboration et al. (2018a) report clear similarities with the bi-modality in the metallicity distribution of the Galactic GCs, which is known to arise due to contributions from clusters predominantly accreted from dwarf galaxies and those formed in situ in the Milky Way. They speculate that the two halo populations must have followed distinct formation channels and surmise that there may be a connection between the two halo HRD sequences and the chemical sequences highlighted earlier in Nissen and Schuster (2010) and Bonaca et al. (2017). Haywood et al. (2018) take a deeper look at the Gaia DR2 HRD to solve the puzzle of the dichotomy of the stellar halo. They demonstrate that while both populations have orbits distinct from the disc, there is a clear difference in the kinematics of the stars in the blue and red sequences. Haywood et al. (2018) show that the red sequence is composed of stars on mostly prograde orbits with lower eccentricities and lower total energies. The blue sequence stars have little angular momentum, have higher eccentricities, and reach out further into the halo. The HRD sequences indeed correspond almost perfectly to the two chemical sequences found by Nissen and Schuster (2010): the red one contains stars with higher $\alpha$-abundances and the blue one with lower $\alpha$-abundances. The connection between kinematics and detailed chemistry is also made in Helmi et al. (2018), who together with Haywood et al. (2018) appear inspired by the earlier analysis of the halo chemistry using APOGEE (see Hawkins et al., 2015; Fernández-Alvar et al., 2018; Hayes et al., 2018). Haywood et al. (2018) associate the red sequence to the thick disc in agreement with Bonaca et al. (2017) and the blue sequence to the ancient massive merger as proposed in e.g. Deason et al. (2013a) and Belokurov et al. (2018b). Haywood et al. (2018) argue that the tightness of the blue sequence in the Gaia HRD is an argument in favour of the single dominant accretion event because contributions from many dwarf progenitors would otherwise bloat it.
@@
Orbital synchronicity. The use of the GCs as tracers highlighted one particular difficulty of the stellar halo studies with Gaia. Put simply, given the limited reach of the Gaia DR2 parallaxes, the vast extent of the halo appeared inaccessible. These observational hindrances can be circumvented using tools of Galactic Dynamics with which one can predict the overall density distribution of the halo using only a small sample of tracer kinematics constrained to the Solar neighourhood (see e.g. Eggen et al., 1962; Chiba and Beers, 2000). This is because the spatial appearance of the Galaxy is nothing but the superposition of the orbits of the individual stars and the density in the configuration space is linked to the kinematics through the continuity of the phase-space flow (Binney and Tremaine, 2008). In a similar vein but in a much more direct experiment, Deason et al. (2018) compute orbits of the metal-rich halo stars in the Solar vicinity and demonstrate that their apo-centres pile-up around Galactocentric distance $r\approx 20$ kpc (see Fig. 8). The location of this pile-up matches the break radius in the stellar halo density (see Section 1.1.2). As a sanity check, they also compare apo-centres of nearby main sequence stars to the apo-centres of the distant BHB stars that were originally used to detect the break and show that stars in the two populations, distant and local, turn around on their orbits at approximately the same distance. Thus three different tracer populations have been shown to have the same orbital behaviour: the nearby main sequence stars, the distant BHBs, and the GCs. Deason et al. (2018) demonstrate in a very candid way that the local metal-rich and radially anisotropic population is only a glimpse of a vast and dominant cloud of stellar debris that dictates the global properties of the stellar halo (e.g. the density break). They also provide the first constraint on the metallicity distribution function of the high-eccentricity population and show that its incidence drops by a factor of $\sim 2$ going from [Fe/H]$=-1$ to [Fe/H]$=-2$.
@@
Helmi et al. (2018) show that on the HRD, these stars are largely confined by isochrones with $-1.3<$[Fe/H]$<-0.9$ and ages of 10-13 Gyr. Helmi et al. (2018) cross-match their selected sample to APOGEE and see two narrow sequences in the $\alpha$-[Fe/H] plane, one mostly flat with $-1<$[Fe/H]$<-0.3$ and [$\alpha$/Fe]$\approx 0.25$ and one sloping down from [Fe/H]$\approx-1.3$, [$\alpha$/Fe] $\approx 0.25$ to [Fe/H]$\approx-0.7$, [$\alpha$/Fe] $\approx 0.1$. Following Bonaca et al. (2017), they assign the high-$\alpha$ sequence to the in-situ population (see also Nissen and Schuster, 2010). Given that the two sequences co-exist over a range of metallicities, Helmi et al. (2018) conclude that the halo stars in the lower $\alpha$-[Fe/H] sequence must have formed in a separate, smaller galaxy. In view of the large spread of metallicity and the presence of low-$\alpha$ high-[Fe/H] stars this dwarf could not be much smaller than the LMC. Their interpretation of the two chemical sequences builds on the analysis of the halo stars in APOGEE by Hawkins et al. (2015), Fernández-Alvar et al. (2018) and Hayes et al. (2018). These earlier works provide a detailed chemo-dynamical comparison of the high- and low-$\alpha$ sequences and agree on the in-situ nature of the former. Crucially, however, in the absence of the Gaia data, these authors refrain from making a claim as to the exact nature of the accreted population, although Fernández-Alvar et al. (2018) report a significantly lower star-formation rate for the low-$\alpha$ population. Helmi et al. (2018)’s conclusion is not solely based on the accumulation of the observational evidence available at the time. Their insight is also the result of an earlier numerical exploration, a series of experiments simulating high mass-ratio mergers between the Milky Way and a dwarf satellite (Villalobos and Helmi, 2008). Helmi et al. (2018) find striking similarities between the Gaia DR2 data and one of the pre-existing simulations (once the simulation is appropriately re-scaled). Helmi et al. (2018) propose to give a name to the progenitor system (and presumably its tidal debris): Gaia-Enceladus.
Following the two papers (Belokurov et al. 2018b and Helmi et al. 2018) that first identified the metal-rich highly-eccentric accreted halo population and proposed that it must be produced in a single massive accretion event, the vast tidal debris cloud is referred to as Gaia Sausage-Enceladus (GS/E).